OpenStudy (anonymous):

Indicate the real part, imaginary part and its conjugate: 4+3\sqrt{-25} / 20

OpenStudy (anonymous):

\[4+3\sqrt{-25}/20\]

OpenStudy (anonymous):

\[\frac{4+3\sqrt{-25}}{20}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or \[4+\frac{3\sqrt{25}}{20}\]

OpenStudy (anonymous):

the first one you got

OpenStudy (anonymous):

\[\sqrt{-25}=\sqrt{25}\sqrt{-1}=5i\]

OpenStudy (anonymous):

so \[\frac{4+3\sqrt{-25}}{20}=\frac{4}{20}+\frac{15}{20}i\]

OpenStudy (anonymous):

or if you care to reduce the fractions ' \[\frac{1}{5}+\frac{3}{4}i\]

OpenStudy (anonymous):

thats the conjugate?

OpenStudy (anonymous):

no, that is the original number

OpenStudy (anonymous):

oh ok, so how would I find the conjugate?

OpenStudy (anonymous):

first of all you have to say what the real part is, and what the imaginary part is do you know it?

OpenStudy (anonymous):

the imaginary part is sqrt-25 , wait do I define these using the answer you got or just the original equation?

OpenStudy (anonymous):

i see you might be confused here the way to write a complex number in standard form is \(a+bi\) not that goofy way your original number was if we do that we get \[\frac{1}{5}+\frac{3}{4}i\]

OpenStudy (anonymous):

that is the number you need to be looking at to answer all the questions

OpenStudy (anonymous):

oh ok so the imaginary is 3/4 and the real is 1/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the conjugate would be 1/5 - 3/4i?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oohh I get it now thank you! :)